a)     A simple closed economy, with an mpc equal to 0.75. The government has passed a balanced budget amendment. The economy goes into a recession, so the government increases government spending by 40 million to try to expand the economy

                   i.           Find the change in output from the increase in government spending.


1
Expert's answer
2021-07-05T17:34:38-0400

the income/spending multiplier = 1/(1-0.75)

= 4 ∆Y = mult × ∆G

∆Y = 4 × 40 million ∆Y = 160 million